From an elevated point A, a stone is projected vertically upward. When the stone reaches a distance h below A, its velocity is double what it was at height h below A. Show that the greatest height obtained by the stone above A is 5h/3.


1
Expert's answer
2021-02-16T06:11:04-0500

Let u be the velocity with which the stone is projected vertically upwards.

"v_{-h}=2V_h \\

v_{-h}^2=4V_h^2 \\

u^2 - 2g - (-h) = 4(u^2 -2gh) \\

u^2 = \frac{10gh}{3} \\

h_{max} = \frac{u^2}{2g}= \frac{5h}{3}"